Suppose test scores on a certain exam follow a normal distribution with μ = 70. If 8% of students scored above 80, what is the standard deviation of the distribution?
Answer #1

Answer:

Given,

Mean μ = 70

x = 80

P(Z > z) = 0.08

from the standard normal distribution table,

P(Z > 1.41) = 0.08

So z = 1.41

we know,

z = (x - μ)/s

substitute known values

1.41 = (80 - 70)/s

s = 10/1.41

s = 7.09

Standard deviation = 7.09